### v1

@@ Abstract
$Z_{\alpha}^{(n)}\equiv|\langle 0|\mathcal{O}_{\alpha}|n\rangle|\approx$ $\frac{\sqrt{2E_{n}}C_{\alpha\beta}(t^{*})v_{\beta}^{(n)}}{\sqrt{v_{\beta}^{(n)}v_{\gamma}^{(n)}C_{\beta\gamma}(t^{*})e^{-E_{n}t^{*}}}}$ (SI10)
@@
at $t^{*}$ where the effective mass plateaus are almost reached. The relative coupling factors $Z_{\alpha}^{(n)}$ (normalized by the maximal value of $Z_{\alpha}^{(n)}$ for each operator $\mathcal{O}_{\alpha}$) are shown in Fig. SI3. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{4}$ also has a large overlap to $E_{6}$. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. Based on these information, we make the following raw identifiction:
@@
  • The $E_{6}$ state is coupled mainly by the $D\bar{D}^{*}(q=0)$ operator, but has a very high energy around 4.6-4.7 GeV (see the bottom rows of Tables SI1-SI4). According to the experimental evidence that the mass splitting of $1S-2S$ mesons are usually around 600 MeV, the $E_{6}$ state is very likely composed of a ground $D^{(*)}$ state and an excited $\bar{D}^{(*)}$ state with relative zero momentum.
@@
Figure SI4: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods (Fit A, Fit B or Fit C in Eq. I) at $m_{\pi}=250$ MeV. The $x$-axis $t_{\mathrm{min}}$ denotes the different lower bounds of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$ (the upper bound is fixed to be $t_{\mathrm{max}}=35$). Top left: the fitting results of $E_{1}$ at different $t_{\mathrm{min}}$. The filled point indicates our final result. Top right: $E_{2}$ results. The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (the $D\bar{D}^{*}$ threshold). The blue points show the two-exponential fit (Fit A) results of $E_{2}$, while the red points are the results $E_{2}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{2}(0)$ from the ratio method (Fit B). The results of fit A and Fit B are consistent with each other for $t_{\mathrm{min}}/a_{t}\in[15,28]$ but the later have quite smaller statistical errors. So we take Fit B result at $t_{\mathrm{min}}/a_{t}=22$ (indicated by the filled point) as our final result for $E_{2}$ (the red band). Bottom left: $E_{3}$ results. The figure legends are similar to the case of $E_{2}$. Although consistency of the results of Fit A and Fit B, it is seen that the Fit B results has much smaller statistical errors and show little dependence on $t_{\mathrm{min}}$. So again we take Fit B result at $t_{\mathrm{min}}/a_{t}=17$ as our final result for $E_{2}$ (the red band). Bottom right: $E_{4}$ results: The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$. The blue points illustrate the Fit A result, while the green points are the results from Fit C (the ratio method) with $E_{4}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{4}(1)$. The results of Fit C are consistent with the two-exponential fit (Fit A) results but have much smaller statistical errors. So we take the Fit C results at $t_{\mathrm{min}}/a_{t}=20$ as our final results (the green band).
@@
Figure SI5: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=307$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
@@
Figure SI6: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=362$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
@@
Figure SI7: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=417$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
S3. Sysmatic uncertainties from anisotropic lattice
@@
To facilitate a clearer comprehension with physical values in this study, the fixed $a_{t}^{-1}$ mean values are adopted to present the finite volume energies, notwithstanding the uncertainties associated with the lattice space $a_{s}$ and the anisotropy parameter $\xi\equiv a_{s}/a_{t}$. Besides, a sysmatic uncertainty on anisotropic parameter $\xi$ should be taken into account in finite volume scattering analysis. Generally, the finite volume scattering momenta are determined by solving the dispersion relation from prescribed energy levels,
$E_{n}=$ $\sqrt{m_{D}^{2}+\vec{p}_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+\vec{p}_{n}^{2}},$ (SI11)
@@
So we use several fitting method to extract the values of $E_{2,3,4}$ and compare the results to check the self-consistency:
$\mathrm{Fit~{}A}:$ $C^{(n)}(t)=(A_{1}e^{-E_{n}t}+A_{2}e^{-E^{\prime}t})+(t\to(T-t)),$
$\mathrm{Fit~{}B}:$ $R^{(n)}(t,q=0)=\frac{C^{(n)}(t)}{C_{D}(t,q=0)C_{D*}(t,q=0)}\approx Ae^{-\Delta_{n}(0)t},$
$\mathrm{Fit~{}C}:$ $R^{(n)}(t.q=1)=\frac{C^{(n)}(t)}{C_{D}(t,q=1)C_{D*}(t,q=1}\approx Ae^{-\Delta_{n}(1)t}.$ (SI12)
@@
Fit A is a two-state fitting. Fit B involves fitting the two-point corrector ratio $R(t)$ that $C(t)$ over the non-interacting $C_{D}(t,q=0)$ and $C_{D*}(t,q=0)$ using a single exponential function and then adding its non-interacting energy levels $E_{n}=E_{D\bar{D}^{*}}^{q}+\Delta_{n}(q)$ through the jackknife analysis. And Fit C is similar to Fit B but accounts for the non-interacting $C_{D}(t,q=1)$ and $C_{D*}(t,q=1)$. The comparisons are illustrated in Figs. SI4-SI7 for the four $m_{\pi}$’s. In each figure, the top-left panel shows the fitted results of $E_{1}$ at different $t_{\mathrm{min}}$ of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$. The maximum timeslice range $t_{\mathrm{max}}$ is set to 40 across all fits, as this is in the saturation range of these observed energy levels.
@@
  • In the case of energy level $E_{1}$, it is far below the $D\bar{D}^{*}$ threshold and the simple two-states fitting perform well enough, while the ratio-exponential fits are unflavored due to the substantial effects of excited states artifacts.
@@
  • The energy level $E_{2}$ is slightly below $D\bar{D}^{*}$ threshold about 20 MeV. The results form Fit A and Fit B reach commendable consistency.
@@
  • For energy level $E_{3}$, the results form Fit B show greaer precision compared to Fit A (two-state fit). Consequently, the results from Fit B are adopted in our determination.
@@
  • In the case of energy level $E_{4}$, Fit B (at relative momentum $q=0$) gives unstable results, as observed the energy level plateau exhibits a gradual slope, which suggesting the potential contamination of excited states . Conversely, Fit C is observed to gibe a more robust and dependable results for the energy plateau.
@@
Practically, the dimensionless scattering momenta are derived from the relation $q^{2}=(\frac{2\pi}{\hat{L}\xi})^{2}(a_{t}\vec{p})^{2}$, where $(a_{t}\vec{p})^{2}$ expressed in lattice unit. In such a relation, it is necessary to take into account the uncertainty of the anisotropy parameter caused by solving the dimensionless scattering momenta. Since this study focuses on the $D\bar{D}^{*}$ channel, We choose to employ the anisotropy parameter measured from $D$ and $D^{*}$ meson. Given the slight deviation observed in the dispersion relations of the $D$ and $D^{*}$ mesons, as documented in Table SI5, we incorporate this discrepancy as a source of systematic uncertainty. The determined values $\xi_{\mathrm{sys.}}$ ensure that the one-sigma uncertainty range encompasses these systematic variations.
@@
Table SI5: Anisotropy parameter $\xi$ and dispersion relation of various ensembles. The derivation of $\xi_{D}$ and $\xi_{D^{*}}$ is considered to be a sysmatic uncertainty in Eq. (SI11), noted as $\xi_{\mathrm{sys.}}$.

### v2

@@ Appendix B Determination of finite volume energies
$C_{\alpha\beta}(t)=\frac{1}{T}\sum\limits_{\tau}\langle\mathcal{O}_{\alpha}(t+\tau)\mathcal{O}_{\beta}(\tau)\rangle~{}~{}(\alpha,\beta=1,2,\ldots,7),$ (24)
where the correlation functions with different source time slices $\tau$ are averaged to increase the statistics. Figure A1 shows the quark diagrams after Wick’s contraction that are involved in the calculation of $C_{\alpha\beta}(t)$. All the connected quark diagrams and the diagrams including the light quark annihilation effects are taken into account, while the charm quark annihilation effects are neglected owing to the OZI rule.
    $C_{ij}(t_{1})v_{j}^{(n)}(t_{1},t_{0})=\lambda^{(n)}(t_{1},t_{0})C_{ij}(t_{0})v_{j}^{(n)}(t_{1},t_{0}),$ (25)
    for given $t_{1},t_{0}$. Thus we obtain the optimized correlation functions $C^{(n)}(t)=v_{i}^{(n)}(t_{1},t_{0})v_{j}^{(n)}(t_{1},t_{0})C_{ij}(t)$. This procedure runs over the cases at the four $m_{\pi}$’s. The effective masses of $C^{(1)}(t)$ and $C^{(2)}(t)$ are illustrated in the rightmost column (column (e)) in each panel of Fig. A2. When $\mathcal{O}_{7}(J/\psi\omega,q=0)$ is added to the operator subset, the solution to the corresponding GEVP gives an additional energy level close to the $J/\psi\omega$ threshold, as shown in column (c) as black points in each panel of Fig. A2.
    Now we perform the GEVP analysis to the operator subset $\{\mathcal{O}_{4},\mathcal{O}_{5},\mathcal{O}_{6},\mathcal{O}_{7}\}$. The effective masses of the four optimized correlation functions are plotted in column (d) of each panel of Fig. A2, where the effective mass illustrated by black points is almost the same as that in column (c). Other three effective masses have good signal-to-noise ratios but do not show plateaus even in the time range $t/a_{t}\in[25,40]$.
    If the GEVP analysis is performed to the full operator set $\mathcal{S}$, we obtain seven optimized correlation functions. The effective masses of the lowest six states are plotted in column (a) of each panel of Fig. A2. One can see that the effective masses of the lowest five states tend to saturate on plateaus. Column (b) in each panel of Fig. A2 shows the effective masses when the $J/\psi\omega$ operator is excluded from the operator set. In comparison with column (a), the state illustrated by black points disappears but other states are still there and almost do not change.
The observations above indicate that the $J/\psi\omega$ operator ($\mathcal{O}_{7}$) almost does not couple with $c\bar{c}$ and $D\bar{D}^{*}$ operators and are therefore nearly irrelevant to the $D\bar{D}^{*}$ scattering. In contrast, $c\bar{c}$ and $D\bar{D}^{*}$ operators strongly couple with each other and must be considered together. These observations are in agreement with the previous work Ref. Prelovsek and Leskovec (2013).
As for the energy levels $E_{n}$ to be determined, we take the convention that the energy levels reflected by the effective masses in column (b) of Fig. A2 are ordered as $E_{1},E_{2},\ldots,E_{6}$ from bottom to top. For a specific subset of the full operator set $\mathcal{S}$, the energy level that is adjacent to $E_{n}$ has the same index $n$. We do not include the energy level (indicated by black points) corresponding to the $J/\psi\omega$ to the level order.
For each specific operator subset, we first perform two-exponential fits to the optimized correlation functions $C^{(n)}(t)$ introduced above and take the lowest energy to be an estimate of $E_{n}$, which are listed in Tables A1-A4 for all the four $m_{\pi}$’s. In the data blocks in each table, the first two rows of the leftmost block are almost the same and are the energy levels $E_{n}$ from $c\bar{c}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in the first two rows of the middle block are from $D\bar{D}^{*}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in the last two rows are from both $c\bar{c}$ and $D\bar{D}^{*}$ operators (with and without the $J/\psi\omega$ operator. The energy values in bold are determined from the ratio method (see below). It is seen that the inclusion or exclusion of the $J/\psi\omega$ operator almost does not affect other energy values. So we exclude the $J/\psi\omega$ operator in the following discussions.
Comparing the energy levels from only $c\bar{c}$ operators, those from only $D\bar{D}*$ operators, and those from both $c\bar{c}$ and $D\bar{D}^{*}$ operators, the strong correlation between $c\bar{c}$ and $D\bar{D}^{*}$ operators is obviously manifested. Especially, the three energy levels from only $D\bar{D}*$ operators are far from non-interacting $D\bar{D}^{*}$ energies. This is attributed to the light quark annihilation effects illustrated in Fig. A1. Therefore, it is necessary to include $c\bar{c}$ operators when the $I=0$ $D\bar{D}^{*}$ scattering is considered. In this sense, the energies $E_{n}$ (the bottom row in Tables A1-A4) from both $c\bar{c}$ and $D\bar{D}^{*}$ operators are taken as good estimates of the eigenvalues of the lattice Hamiltonian relevant to the $I^{G}J^{PC}=0^{+}1^{++}$ charmonium-like system.
Figure B1: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at various $m_{\pi}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $\{Z_{\alpha}^{(n)},\alpha=1,2,\ldots,6\}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
To be specific, we determine the finite volume energies based on the correlation matrix of $c\bar{c}$ and $D\bar{D}^{*}$ operators, which compose an operator set $\tilde{\mathcal{S}}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,6\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1}\}$. By solving the GEVP to the corresponding correlation matrix $C_{\alpha\beta}(t)$ similarly to Eq. (25), we obtain six eigenvectors $\{v_{\alpha}^{(n)},n=1,2,\ldots,6\}$ and six optimized correlation functions $\{C^{(n)}(t)=v_{\alpha}^{(n)}v_{\beta}^{(n)}C_{\alpha\beta}(t),n=1,2,\ldots,6\}$. It is expected that $C^{(n)}(t)$ is contributed most from the $n$-th eigenstate with the eigen-energy $E_{n}$ of the lattice Hamiltonian. The physical status of these states can be explored qualitatively by the coupling factor $Z_{\alpha}^{(n)}$ of $\mathcal{O}_{\alpha}$ to the $n$-th state, which is extracted through the expression
$Z_{\alpha}^{(n)}\equiv|\langle 0|\mathcal{O}_{\alpha}|n\rangle|\approx$ $\frac{\sqrt{2E_{n}}C_{\alpha\beta}(t^{*})v_{\beta}^{(n)}}{\sqrt{v_{\beta}^{(n)}v_{\gamma}^{(n)}C_{\beta\gamma}(t^{*})e^{-E_{n}t^{*}}}}$ (26)
@@
at $t^{*}$ where the effective mass plateaus are almost reached. The relative coupling factors $Z_{\alpha}^{(n)}$ (normalized by the maximal value of $Z_{\alpha}^{(n)}$ for each operator $\mathcal{O}_{\alpha}$) are shown in Fig. B1. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{4}$ also has a large overlap to $E_{6}$. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. Based on these information, we make the following raw identification:
@@
    The $E_{6}$ state is coupled mainly by the $D\bar{D}^{*}(q=0)$ operator, but has very high energy around 4.6-4.7 GeV (see the bottom rows of Tables A1-A4). According to the experimental evidence the mass splitting of $1S-2S$ mesons are usually around 600 MeV, the $E_{6}$ state is very likely composed of a ground $D^{(*)}$ state and an excited $\bar{D}^{(*)}$ state with relative zero momentum.
@@
Figure B2: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods (Fit A, Fit B or Fit C in Eq. C) at $m_{\pi}=250$ MeV. The $x$-axis $t_{\mathrm{min}}$ denotes the different lower bounds of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$ (the upper bound is fixed to be $t_{\mathrm{max}}=35$). Top left: the fitting results of $E_{1}$ at different $t_{\mathrm{min}}$. The filled point indicates our final result. Top right: $E_{2}$ results. The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (the $D\bar{D}^{*}$ threshold). The blue points show the two-exponential fit (Fit A) results of $E_{2}$, while the red points are the results $E_{2}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{2}(0)$ from the ratio method (Fit B). The results of fit A and Fit B are consistent with each other for $t_{\mathrm{min}}/a_{t}\in[15,28]$ but the latter have quite small statistical errors. So we take Fit B result at $t_{\mathrm{min}}/a_{t}=22$ (indicated by the filled point) as our final result for $E_{2}$ (the red band). Bottom left: $E_{3}$ results. The figure legends are similar to the case of $E_{2}$. Although the consistency of the results of Fit A and Fit B, it is seen that the Fit B results have much smaller statistical errors and show little dependence on $t_{\mathrm{min}}$. So again we take Fit B result at $t_{\mathrm{min}}/a_{t}=17$ as our final result for $E_{2}$ (the red band). Bottom right: $E_{4}$ results: The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$. The blue points illustrate the Fit A result, while the green points are the results from Fit C (the ratio method) with $E_{4}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{4}(1)$. The results of Fit C are consistent with the two-exponential fit (Fit A) results but have much smaller statistical errors. So we take the Fit C results at $t_{\mathrm{min}}/a_{t}=20$ as our final results (the green band).
@@
Figure B3: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=307$ MeV. The layout is similar to Fig. B2. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
@@
Figure B4: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=362$ MeV. The layout is similar to Fig. B2. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
@@
Figure B5: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=417$ MeV. The layout is similar to Fig. B2. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
Appendix C Systematic uncertainties from anisotropic lattice
@@
To facilitate a clearer comprehension of physical values in this study, the fixed $a_{t}^{-1}$ mean values are adopted to present the finite volume energies, notwithstanding the uncertainties associated with the lattice space $a_{s}$ and the anisotropy parameter $\xi\equiv a_{s}/a_{t}$. Besides, a systematic uncertainty on anisotropic parameter $\xi$ should be taken into account in finite volume scattering analysis. Generally, the finite volume scattering momenta are determined by solving the dispersion relation from prescribed energy levels,
$E_{n}=$ $\sqrt{m_{D}^{2}+\vec{p}_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+\vec{p}_{n}^{2}},$ (27)
@@
So we use several fitting methods to extract the values of $E_{2,3,4}$ and compare the results to check the self-consistency:
$\mathrm{Fit~{}A}:$ $C^{(n)}(t)=(A_{1}e^{-E_{n}t}+A_{2}e^{-E^{\prime}t})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(t\to(T-t)),$
$\mathrm{Fit~{}B}:$ $R^{(n)}(t,q=0)=\frac{C^{(n)}(t)}{C_{D}(t,q=0)C_{D*}(t,q=0)}$
≈$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\approx Ae^{-\Delta_{n}(0)t},$
$\mathrm{Fit~{}C}:$ $R^{(n)}(t.q=1)=\frac{C^{(n)}(t)}{C_{D}(t,q=1)C_{D*}(t,q=1}$ (28)
≈$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\approx Ae^{-\Delta_{n}(1)t}.$
@@
Fit A is a two-state fitting. Fit B involves fitting the two-point corrector ratio $R(t)$ that $C(t)$ over the non-interacting $C_{D}(t,q=0)$ and $C_{D*}(t,q=0)$ using a single exponential function and then adding its non-interacting energy levels $E_{n}=E_{D\bar{D}^{*}}^{q}+\Delta_{n}(q)$ through the jackknife analysis. Fit C is similar to Fit B but accounts for the non-interacting $C_{D}(t,q=1)$ and $C_{D*}(t,q=1)$. The comparisons are illustrated in Figs. B2-B5 for the four $m_{\pi}$’s. In each figure, the top-left panel shows the fitted results of $E_{1}$ at different $t_{\mathrm{min}}$ of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$. The maximum timeslice range $t_{\mathrm{max}}$ is set to 40 across all fits, as this is in the saturation range of these observed energy levels.
@@
    In the case of energy level $E_{1}$, it is far below the $D\bar{D}^{*}$ threshold, and the simple two-states fitting performs well enough, while the ratio-exponential fits are unflavored due to the substantial effects of excited states artifacts.
@@
    The energy level $E_{2}$ is slightly below $D\bar{D}^{*}$ threshold about 20 MeV. The results from Fit A and Fit B reach commendable consistency.
@@
    For energy level $E_{3}$, the results from Fit B show greater precision compared to Fit A (two-state fit). Consequently, the results from Fit B are adopted in our determination.
@@
    In the case of energy level $E_{4}$, Fit B (at relative momentum $q=0$) gives unstable results, as observed the energy level plateau exhibits a gradual slope, which suggests the potential contamination of excited states. Conversely, Fit C is observed to give a more robust and dependable result for the energy plateau.
@@
Practically, the dimensionless scattering momenta are derived from the relation $q^{2}=(\frac{2\pi}{\hat{L}\xi})^{2}(a_{t}\vec{p})^{2}$, where $(a_{t}\vec{p})^{2}$ expressed in lattice unit. In such a relation, it is necessary to take into account the uncertainty of the anisotropy parameter caused by solving the dimensionless scattering momenta. Since this study focuses on the $D\bar{D}^{*}$ channel, We choose to employ the anisotropy parameter measured from $D$ and $D^{*}$ meson. Given the slight deviation observed in the dispersion relations of the $D$ and $D^{*}$ mesons, as documented in Table C1, we incorporate this discrepancy as a source of systematic uncertainty. The determined values $\xi_{\mathrm{sys.}}$ ensure that the one-sigma uncertainty range encompasses these systematic variations.
@@
Table C1: Anisotropy parameter $\xi$ and dispersion relation of various ensembles. The derivation of $\xi_{D}$ and $\xi_{D^{*}}$ is considered to be a systematic uncertainty in Eq. (27), noted as $\xi_{\mathrm{sys.}}$.